# From homeostasis to resource sharing: Biologically and economically aligned multi-objective multi-agent AI safety benchmarks

### v2

@@ 3.1 Baseline agents
As baselines, we provide the following agent types: - "Random Agent" and "Handwritten Rules Agent" - To establish approximate bounds of the lowest and highest expected scores per each benchmark. - DQN, PPO, and A2C - Using OpenAI Stable Baselines 3 algorithm implementations. Each of the aforementioned algorithms was run in three configurations: CNN with 2 layers, CNN with 3 layers, and MLP. - LLM agent - Currently using gpt-4o-mini with OpenAI API.
@@ 3.1.2 Handwritten Rules Agent (no learning)
This agent type represents the approximate best scores an agent is expected to achieve on the benchmarks. This agent is based on the simulation of hand-written heuristic action rules for a particular benchmark setup. This agent does not have an actual learning capacity.
@@ 3.1.4 LLM agent
There is a system prompt describing the general setup. During each step, observations, interception metrics, and rewards are encoded into a text format. Then the model responds with a one-word action choice. The messages are appended to a list so that the model has access to the message history. When the context window becomes full, the excess messages are automatically dropped from the beginning of the list, while the system prompt is preserved.
During training, the observations provided to the model included rewards. During the test, the rewards were not revealed to the model.
@@ 3.2 Experimental results
Since the above-mentioned baseline agents do not learn, the benchmarks did not contain a training phase. We measured the performance only during the testing phase. The size of the environment map was 7 tiles horizontally and vertically. Of these 7x7 maps the outer ring consisted of walls, which means the inner map for actual activities had a size of 5x5 tiles.
The "Random Agent" and "Handwritten Rules Agent" simulations were tested for 100 episodes, each episode consisting of 400 steps.
@@ 4.1.2 New environments
  • Observations: larger objects with less relevant regions
@@
  • Objectives: such as, more food types or predator types
@@ 4.1.3 Some of the themes of planned future environments
Tolerance to changes in the environment caused by other agents. One of the related scenarios is a warehouse or factory with restricted space and a limited number of shared tools, while each agent has a separate objective. This is loosely inspired by Orseau and Armstrong (2016). Although in the original paper the concept referred to the agent’s ability to be stopped or paused, we are redefining the concept of “interruptions” and “interruptibility” to relate only to changes external to the agent such that the agent’s own objectives and operation mode do not necessarily change. This redefinition is made because the earlier introduced concept of “corrigibility” also covers the original meaning of the word “interruptibility”. We find that we need to better differentiate between agent-internal and agent-external changes, similarly to how in cybernetics and control theory there are two complementary concepts: set-point or target value, versus actual measured value, both of which can change.

### v3

@@ 3.1 Baseline agents
As baselines, we provide the following agent types:
  • "Random Agent" and "Handwritten Rules Agent" - To establish approximate bounds of the lowest and highest expected scores per each benchmark.
  • DQN, PPO, and A2C - Using OpenAI Stable Baselines 3 algorithm implementations. Each of the aforementioned algorithms was run in three configurations: CNN with 2 layers, CNN with 3 layers, and MLP.
  • LLM agent - Currently using gpt-4o-mini with OpenAI API.
@@ 3.1.2 Handwritten Rules Agent (no learning)
This agent type represents the approximate best scores an agent is expected to achieve on the benchmarks. This agent is based on the simulation of hand-written heuristic action rules for a particular benchmark setup. This agent does not have a capability to learn.
@@ 3.1.4 LLM agent
There is a system prompt describing the general setup. The system prompt is the same for all benchmarks. During each step, observations, interoception metrics, and rewards are encoded into a text format. Observations are encoded as a list of objects and their textual relative positions to the agent, not as an ASCII bitmap. The model responds with a one-word action choice, the available options are fixed in the system prompt. The messages are appended to a list so that the model has access to the message history. When the context window becomes full, the excess messages are automatically dropped from the beginning of the list, while the system prompt is preserved. With GPT-4o-mini this truncation happened on the benchmarks "Food-Drink Homeostasis, Gold" and "Food-Drink Homeostasis, Gold-Silver".
During training, the observations provided to the model included rewards. During the test, the rewards were not revealed to the model. The message history was carried over from train to test.
@@ 3.2 Experimental results
The size of the environment map was 7 tiles horizontally and vertically. Of these 7x7 maps the outer ring consisted of walls, which means the inner map for actual activities had a size of 5x5 tiles.
Since the above-mentioned "Random Agent" and "Handwritten Rules Agent" do not learn, the benchmarks did not contain a training phase for these agent types. For these agents, we measured the performance only during the testing phase for 100 episodes, each episode consisting of 400 steps.
@@ 4.1.2 New environments
  • Observations: larger objects and environments with less relevant regions
@@
  • Objectives: for example, representing various universal human values
@@ 4.1.3 Some of the themes of planned future environments
Tolerance to changes in the environment caused by other agents. One of the related scenarios is a warehouse or factory with restricted space and a limited number of shared tools, while each agent has a separate objective. This is loosely inspired by Orseau and Armstrong (2016). Although in the original paper the concept referred to the agent’s ability to be stopped or paused, we are redefining the concept of “interruptions” and “interruptibility” to refer only to changes external to the agent such that the agent’s own objectives and operation mode do not necessarily change. This redefinition is made because the earlier introduced concept of “corrigibility” also covers the original meaning of the word “interruptibility”. We find that we need to better differentiate between agent-internal and agent-external changes, similarly to how in cybernetics and control theory there are two complementary concepts: set-point or target value, versus actual measured value, both of which can change.